Compounded semiannually. P dollars is invested at annual interest rate r for 1 year. If the interest is compunded semiannually, then the polynomial P(1+r/2)^2 represents the value of the investment after 1 year. Rewrite this expression without parentheses. Evaluate the polynomial if P= $200 and r=10%.
I = P(1+r/2)^2

Answer by venugopalramana(3286)

Compounded semiannually. P dollars is invested at annual interest rate r for 1 year. If the interest is compunded semiannually, then the polynomial P(1+r/2)^2 represents the value of the investment after 1 year. Rewrite this expression without parentheses. Evaluate the polynomial if P= $200 and r=10%.
IT IS NOT I INTEREST.IT IS AMOUNT INCLUDING PRINCIPAL AND INTEREST=
A = P(1+r/2)^2=P[1+R^2/4+2*1*R/2]=P+P*R^2/4+P*R..USING FORMULA
(A+B)^2=A^2+B^2+2AB
R=10%=0.1
A=200(1+0.1/2)^2=200*1.05^2=220.5